A monopolist faces two totally separated markets with inverse demand p=100 – qA and p=160−2qB respectively. The monopolist has no fixed costs and a marginal cost given by mc= 2 /3 q Find the profit maximizing total output and how much of it that is sold on market A and market B respectively if the monopoly uses third degree price discrimination.
a) What prices will our monopolist charge in the two separate markets?
1
Expert's answer
2020-06-08T11:06:14-0400

a

For monopoly must be "MC=p"


For firm A

"\frac{2}{3}Q=100-Q_A"


"\frac {4}{3}Q_A=100"


"Q_A=75"


"p=25"

For firm B



"\frac{2}{3}Q_B=160-2Q_B"


"\frac {8}{3}Q_B=160"

"Q_B=60"


"p=100"

b



"Q_A=100-p"


"Q_A^\/=-1"


"E_A=Q_A^/\frac{p_A}{Q_A}"


"E_A=-\frac {25}{75}=-\frac{1}{3}"


"Q_B=80-0.5p"


"Q_B^\/=-0.5"


"E_B=-0.5 \frac{60}{100}=-0.3"

The absolute values ​​of elasticity at the points characteristic for equilibrium in these markets are mutually inverse to each other.